2 stones A & B are dropped from 2 different towers such that they travel 44.1m and 63.7m in their last second of their motion respectively. find the ratio of the height of the 2 towers from where the stones are dropped?????????????????
plzzzzzz answer

Grade:9

1 Answers

Sanjeev Malik IIT BHU
12 years ago

take stone 1 total time = x                                                    """" ratio h1/h2= x2/y2   """""

------------2------------= y   initial speed 0 for both

dis in last sec for 1= 1/2* g*(x2-(x-1)2)=1/2*g* (2x-1)= 44.1

--------------------2= 1/2*g* (y2-(y-1)2)=1/2*g* (2y-1)= 63.7

thus find x and y easily

and put the value to get ratio of heights.

 

if helpful approve by clicking "yes" below